OpenStudy (erinweeks):

HELP PLEASE (: What is the remainder of (x3 – 6x – 9x + 3) ÷ (x - 3) A. -51 B. -51 x - 3 C. -17 x D. -17 x - 3

OpenStudy (compassionate):

(-x)-6x+(-3)3 ------------- = - 7x x - 3

OpenStudy (erinweeks):

there is no -7x ?

OpenStudy (compassionate):

What form is it in?

OpenStudy (erinweeks):

it doesnt say ? what posted was all it said .

OpenStudy (erinweeks):

A. -51 B. \[\frac{ -51 }{ x-3 }\] C. \[\frac{ -17 }{ x }\] D.\[\frac{ -17 }{ x-3 }\]

OpenStudy (erinweeks):

them are the answers i can choose from

jimthompson5910 (jim_thompson5910):

Use the remainder theorem

jimthompson5910 (jim_thompson5910):

If you divide P(x) by x-k, then the remainder 'r' is r = P(k)

OpenStudy (compassionate):

Sorry, I was solving and not giving the remainder.

OpenStudy (erinweeks):

do you guys no what the remaidder is ?

jimthompson5910 (jim_thompson5910):

yes, what do you get for the remainder?

OpenStudy (erinweeks):

well i thought it was a but i got it wrong ..

OpenStudy (erinweeks):

-51

jimthompson5910 (jim_thompson5910):

It is -51, here's why x^3 - 6x^2 - 9x + 3 (3)^3 - 6(3)^2 - 9(3) + 3 -51

jimthompson5910 (jim_thompson5910):

Now because the remainder is a part of the divisor (basically its what is left over), you can write "remainder -51" as -51/(x-3)

OpenStudy (erinweeks):

but its not because i got it wrong on my test ..

OpenStudy (erinweeks):

So Could B , be the answer ?

jimthompson5910 (jim_thompson5910):

yes, if A doesn't work, then it has to be B

OpenStudy (erinweeks):

okay thank you so much !

jimthompson5910 (jim_thompson5910):

you're welcome